Using the simply money multiplier model, what quantity of securities must the Federal Reserve purchase to generate an increase in the size of checkable deposits by $22,500, assuming the required reserve ratio is 4%?

Group of answer choices

A) 810

B) 850

C) 900

D) 920
Answer #1

multiplier = 1/r = 1/0.04 = 25

So, to increase the size of checkable deposits by 22500, the quantity of securities that the Fed needs to purchase

= 22500/25

= 900

option(C)